# Neutrino oscillations originate from virtual excitation of Z bosons

### v1

@@ Abstract
To account for neutrino oscillations, it is postulated that the neutrino has nonvanishing mass and each flavor eigenstate is formed by three distinct mass eigenstates, whose probability amplitudes interfere with each other during its propagation. However, I find that the energy conservation law requires these mass eigenstates, if they exist, to be entangled with distinct joint energy eigenstates of the other particles produced by the same weak interaction as the neutrino. This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which are responsible for flavor oscillations under the aforementioned postulation. I reveal that the neutrino oscillations actually originate from virtual excitation of the Z bosonic field diffusing over the space. During the propagation, the neutrino can continually excite and then immediately re-absorb a virtual Z boson. This virtual bosonic excitation produces a backaction on the neutrino, enabling it to oscillate among three flavors. When the neutrino propagates in matter, its behavior is determined by the competition between the coherent flavor transformation and decoherence effect resulting from scatterings.
@@
Suppose that the entire system just before participating in the weak interaction (e.g., the neutron just before making a $\beta$ decay) has a definite total energy $E_{T}$. Then the entire state of the neutrino and accompanying particles produced by such an interaction can be written as $\mathop{\displaystyle\sum}\limits_{i}C_{i}\left|\nu_{i}\right\rangle\left|\psi_{i}\right\rangle$. The energy of $\left|\psi_{i}\right\rangle$ should be $E_{i}=E_{T}-\sqrt{p^{2}-m_{i}^{2}}$, where $p$ represents the momentum of the neutrino, and $m_{i}$ is the mass associated with $\left|\nu_{i}\right\rangle$. Only with such an arrangement can the final system has the same definite total energy, $E_{T}$, as the initial one, which ensures the process to comply with the energy conservation law. Under the assumption $m_{1}\neq m_{2}\neq m_{3}$, we have $E_{1}\neq E_{2}\neq E_{3}$, which implies that $\left|\psi_{1}\right\rangle$, $\left|\psi_{2}\right\rangle$, and $\left|\psi_{3}\right\rangle$ correspond to three energy eigenstates with different eigenvalues for the subsystem combined by the accompanying particles, thereby being orthogonal to each other. Due to this orthogonality, when the freedom degrees of the accompanying particles are traced out, the neutrino is left in a classically mixed state, described by the density operator, $\rho=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}\right|^{2}\left|\nu_{i}\right\rangle\left\langle\nu_{i}\right|$. Consequently, there does not exist any quantum coherence among the mass eigenstates, so that their mass differences cannot result in any flavor oscillation.
Moreover, if neutrino flavor eigenstates are defined as superpositions of mass eigenstates, it is impossible to creat a definite flavor by any weak interaction. This is due to the fact that, for the classical mixture $\rho$, the populations of the three flavor eigenstates ($\left|\nu_{\alpha}\right\rangle$) are all nonzero, given by $P_{\alpha}=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}G_{i\alpha}\right|^{2}$, where $\left|\nu_{i}\right\rangle=\mathop{\displaystyle\sum}\limits_{i}G_{i\alpha}\left|\nu_{\alpha}\right\rangle$. Therefore, the postulation that the flavor eigenstates are inconsistent with mass eigenstates not only fails to explain neutrino oscillations, but also leads to the wrong conclusion that the neutrino produced by any weak interaction has mixed flavors.

### v2

@@ Abstract
To account for neutrino oscillations, it is postulated that the neutrino has nonvanishing mass and each flavor eigenstate is formed by three distinct mass eigenstates, whose probability amplitudes interfere with each other during its propagation. However, I find that the energy conservation law requires these mass eigenstates, if they exist, to be entangled with distinct joint energy eigenstates of the other particles produced by the same weak interaction as the neutrino. This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which is responsible for flavor oscillations under the aforementioned postulation. I reveal that the neutrino oscillations actually originate from virtual excitation of the Z bosonic field diffusing over the space. During the propagation, the neutrino can continually excite and then immediately re-absorb a virtual Z boson. This virtual bosonic excitation produces a backaction on the neutrino, enabling it to oscillate among three flavors. When the neutrino propagates in matter, its behavior is determined by the competition between the coherent flavor transformation and decoherence effect resulting from scatterings.
@@
Suppose that the entire system just before participating in the weak interaction (e.g., the neutron just before making a $\beta$ decay) has a definite total energy $E_{T}$. Then the entire state of the neutrino and accompanying particles produced by such an interaction can be written as $\mathop{\displaystyle\sum}\limits_{i}C_{i}\left|\nu_{i}\right\rangle\left|\psi_{i}\right\rangle$. The energy of $\left|\psi_{i}\right\rangle$ should be $E_{i}=E_{T}-\sqrt{p^{2}+m_{i}^{2}}$, where $p$ represents the momentum of the neutrino, and $m_{i}$ is the mass associated with $\left|\nu_{i}\right\rangle$. Only with such an arrangement can the final system has the same definite total energy, $E_{T}$, as the initial one, which ensures the process to comply with the energy conservation law. Under the assumption $m_{1}\neq m_{2}\neq m_{3}$, we have $E_{1}\neq E_{2}\neq E_{3}$, which implies that $\left|\psi_{1}\right\rangle$, $\left|\psi_{2}\right\rangle$, and $\left|\psi_{3}\right\rangle$ correspond to three energy eigenstates with different eigenvalues for the subsystem combined by the accompanying particles, thereby being orthogonal to each other. Due to this orthogonality, when the freedom degrees of the accompanying particles are traced out, the neutrino is left in a classically mixed state, described by the density operator, $\rho=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}\right|^{2}\left|\nu_{i}\right\rangle\left\langle\nu_{i}\right|$. Consequently, there does not exist any quantum coherence among the mass eigenstates, so that their mass differences cannot result in any flavor oscillation.